Yohe Technology’s stock is expected to pay a dividend of $2.00 a share at the end of the year. The stock currently has a price of $40 a share, and the stock’s dividend is expected to grow at a constant rate of g percent a year. The stock has a beta of 1.2. The market risk premium, kM – kRF, is 7 percent and the risk-free rate is 5 percent. What is the expected price of Yohe’s stock 5 years from today?



1
Expert's answer
2021-11-08T19:38:11-0500

Calculate cost of equity using CAPM model:

(k)=Risk free rate+Beta"\times" Market risk free premium


"=" "5\%+1.2 \times7\%"


"=13.4\%"


Calculate growth rate of dividend using Gordon Growth Dividend model:

"P_{o}=\frac{D_{1}}{k-g}"


"\$40=\frac{\$2}{13.4\%-g\%}"


"\$5.36-\$40g=\$2"


"g=8.4\%"


Calculate terminal value at the end of year 5:

"=\frac {D_{1}\times(1+g\%)^5}{k-g}"


"=\frac {\$2\times(1+8.4\%)^5}{13.4\%-8.4\%}"

"=\$59.87"


"\therefore" the expected price of Yohe’s stock 5 years from today is "\$59.87"